OpenStudy (anonymous):

solve for the variable 2c + 12 = 24

OpenStudy (anonymous):

6

OpenStudy (anonymous):

your pretty fast

OpenStudy (anonymous):

thanks xD

OpenStudy (anonymous):

i just wanted to give a medal

jimthompson5910 (jim_thompson5910):

2c + 12 = 24 2c + 12 - 12 = 24 - 12 2c = 12 2c/2 = 12/2 c = 6

OpenStudy (anonymous):

6